[figure: 9.7 cm]

Spin triplet superconductivity in Sr_{2}RuO_{4}

Karol I. Wysokiński^{1}, G. Litak^{2}, J.F. Annett^{3}, B.L. Györffy^{3} 1Institute of Physics, M. Curie – Skłodowska University, ul. Radziszewskiego 10,

20-031 Lublin, Poland

2Department of Mechanics, Technical University of Lublin, Nadbystrzycka 36, 20-618

Lublin, Poland

3H.H. Wills Physics Laboratory, University of Bristol, Tyndall Ave, Bristol, BS8 1TL, UK

Sr_{2}RuO_{4} is at present the best candidate of the solid state superconducting analogue of the triplet superfluidity in ^{3}He. The material is a good (albeit correlated) Fermi liquid in the normal state and an exotic superconductor below T_{c}. The mechanism of superconductivity and symmetry of the order parameter remain main puzzling issues of on-going research. In the talk I shall present the results of our search for the correct superconducting state realised in the material. Our calculations are based on a threedimensional three-band effective model with a realistic band structure. We have found a state with non-zero order parameter at each of the three sheets of the Fermi surface. The corresponding gap in the quasiparticle spectrum has line or point nodes on the α and β sheets and is complex with no nodes at the γ sheet. This state describes remarkably well a number of existing experiments including power low temperature dependence of the specific heat, penetration depth, thermal conductivity etc. The stability of the state with respect to disorder and spin-orbit coupling will be also briefly analysed.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : invited talk

Corresponding author : Karol I. Wysokiński

Address for correspondence : Institute of Physics, M. Curie – Skłodowska University, ul. Radziszewskiego 10, 20-031 Lublin, Poland

Email address : karol@tytan.umcs.lublin.pl